# Robustifying Long-term Human-Robot Collaboration through a Hierarchical and Multimodal Framework

### v1

@@ IV-C1 Action Prediction
Figure 5 illustrates our action prediction pipeline. Our approach does not simply predict human actions from the current pose trajectory. Instead, we initially predict the future trajectory and subsequently infer the action based on both the current and predicted trajectories. This method ensures that the action prediction is informed by anticipated future movements. Research has demonstrated that joint training of the action and trajectory prediction models enhances performance compared to training standalone action prediction models [2]. We train both the trajectory and action prediction models using supervised learning, utilizing data collected offline. The trajectory prediction model is optimized using regression loss (specifically L2 loss for predicted trajectories), while the action prediction model employs classification loss (namely cross-entropy loss for predicted actions). Once trained, the action prediction model $\hat{f}_{\text{action}}$ applies to predict human actions $\hat{x}_{H,\text{action}}^{t}$ based on the current human pose: $\hat{x}_{H,\text{action}}^{t}=\hat{f}_{\text{action}}(\hat{z}_{H}^{t}).$
@@
In long-term HRC, prediction models designed for one human often fail to generalize across different users due to variations in personal preferences and behavioral styles. To address this, we implement an online adaptation method during deployment [18]. This approach allows the robot to continually refine its predictions by incorporating real-time behavior observations of each user [1]. Our joint training framework facilitates the dynamic adaptation of the action prediction model to various users in real-time HRC scenarios. Specifically, the trajectory prediction model is optimized online using observed trajectory feedback. Since the action prediction depends on the trajectory prediction module, any updates to the trajectory model also enhance action prediction accuracy [22]. For more details on this online adaptation approach in this paper, refer to the work [24]. This technique significantly improves prediction performance, ensuring that the robot’s responses accurately align with each user’s behaviors and intentions.
IV-C2 Plan Prediction
@@
After obtaining predicted action $\hat{x}_{H,\text{action}}^{t}$, the robot predicts the high-level intention $\hat{x}_{H}^{t}$, or the human plan. his involves integrating the current task progress $P_{t}$, which encapsulates the sequence of robot actions executed thus far, with the task graph $\mathcal{G}$. The robot identifies the reference action sequence $R^{\ast}$ within the task graph that most closely aligns with the recorded task progress. This alignment is formulated as:
@@
where $d$ measures the distance between the observed task progress $P_{t}$ and each possible reference sequence $R$ within the task graph. The distance metric $d$ is calculated using the dynamic time warping (DTW) algorithm, renowned for its capability to flexibly align sequences with varying lengths and timings [27]. The result, $R^{\ast}$, represents the best-matched reference action sequence. Note that $R^{\ast}$ may not be unique, since the multiple action sequences may present the same distances due to the potential for the current task progress to align with different action sequences leading to the goal. Following this alignment, the human plan prediction is obtained by refining the predicted human action based on the most probable action class within $R^{\ast}$. This adjustment is expressed as:
$\hat{x}_{H}=R^{\ast}(\hat{x}_{H,\text{action}}^{t}),$ (8)
where $\hat{x}_{H}$ represents the predicted human plan. This equation aligns $\hat{x}_{H,\text{action}}^{t}$ with the most likely action within the set of potential action sequences in $R^{\ast}$. This hierarchical framework effectively integrates low-level human action predictions with high-level task plans. By ensuring that the robot’s predictions are consistently aligned with the task graph and progress, it significantly enhances the accuracy and reliability of human-robot collaboration.
IV-D Speech Recognition and Robot Motion Planning
For the audio modality $O_{A}$, the robot recognizes the audio-based human intention or speech command, $\hat{s}_{H}^{t}$, using a pretrained speech recognition model $\hat{f}_{\rm speech}$: $\hat{s}_{H}^{t}=\hat{f}_{\rm speech}(O_{A})$. In deployment, this model’s parameters are fixed. To enhance the quality of speech recognition, a Voice Activity Detector (VAD) is utilized to filter the input. This VAD activates the speech recognition module only when significant audio signal energy, not noise, is detected.
Upon acquiring the vision-based human intention $\hat{x}_{H}^{t}$ and the speech command $\hat{s}_{H}^{t}$, the robot’s motion planner generates the robot motions from both modalities as follows:
$q_{V}^{t:t+m}$ $=M_{R}(\hat{x}_{H}^{t},P^{t}),~{}~{}q_{A}^{t:t+m}=M_{R}(\hat{s}_{H}^{t},P^{t}),$ (9)
$q^{t:t+m}$ $=(1-w^{t})q_{V}^{t:t+m}+w^{t}q_{A}^{t:t+m},$ (10)
where $q_{V}^{t:t+m}$ and $q_{A}^{t:t+m}$ represent the motion sequences planned based on vision-derived intentions and audio commands, respectively. The final robot motion, $q^{t:t+m}$, integrates these motions, weighted by a scalar $w_{t}$. In our work, we implement an indicator function that gives priority to audio-based observations. Specifically, if keywords are recognized in the speech recognition output, we set $w^{t}=1$, engaging the audio-based motion planner exclusively. Otherwise, the vision-based motion planner predominates, and $w^{t}$ is set to zero. The motion controller operates as a feedback controller with an update frequency of approximately 30 Hz.
@@ V-A Toy Car Assembly Task
We evaluate the effectiveness of our proposed framework through a long-term HRC task, specifically the toy car assembly task. This assembly process incorporates various components, including both short and long tubes, connectors, screws, wheels, and panels. This task is divided into four stages: assembling the bottom, the middle, the top of the car, and installing the wheels, as shown in fig. 1. For simplicity, we will refer to the first three stages as The Bottom, The Middle, and The Top. The assembly order for The Middle and The Top is flexible, allowing variations in the sequence order. Each stage is designed to require collaborative efforts between the human and the robot. The actions and motion sequences within these stages are non-deterministic, offering a range of valid action sequences that can successfully complete each sub-task. This flexibility means that the specific sequence of actions may vary according to individual human preferences. In this scenario, the robot’s role is to support the human collaborator by delivering necessary objects and performing specific assistive actions, such as rotating or lifting objects.
As shown in the bottom part of fig. 3, the tools and objects needed for the assembly task are positioned across different locations. Connectors, screws, and other small materials are placed on a table near the human to allow easy access. Conversely, the tubes are stored on a table that is closer to the robot arm and relatively farther from the human. During the assembly process, the robot assists based on the human’s intentions and the current task progress. For instance, if the human intends to reach for connectors, the robot responds by delivering the appropriate tube—either short or long—based on the observed signals. When the human intends to retrieve screws, the robot aids by spinning the base-object to facilitate easier handling. Similarly, when the task involves attaching wheels, the robot assists by lifting the car’s body to better position it for wheel installation. The source code and demonstration video for the HRC toy car assembly task are available at https://github.com/intelligent-control-lab/Robust-Hierarchial-Multimodal-HRC.
@@ V-B1 Implementation Details
  • Human (Object) Detection. We use a pretrained MobileNet model [28] to detect and classify 21 object types from RGB data. For our purposes, we specifically focus on ”human” detections, applying a confidence threshold of 0.6.
  • Human Pose Detection. The pretrained BlazePose model [29] extracts 33 key points, of which we focus on 15 upper-body points for action and intention prediction.
  • Human Action Prediction. We implement the DLinear model [30] for action and trajectory prediction, which decomposes input time series into trend and seasonal components. This model is trained using offline data.
  • Speech Recognition. The DeepSpeech model [31], an open-source Speech-to-Text engine is utilized for recognizing and processing speech commands.
@@ V-B2 Hypothesis
  • H1: The framework enhances the efficiency of HRC.
@@
  • H3: The framework increases task completion rates in collaborative tasks.
@@
  • H5: The framework improves user flexibility and satisfaction in collaboration tasks.
@@ V-B3 User Study
We conducted a user study with 10 participants, focusing on 3 tasks from a toy car assembly problem. The study included three experimental settings: vision-only, audio-only, and multimodal, each repeated 4 times per participant. The tasks involved two assembly stages and two robot actions:
  • Task 1: Pick and place objects based on human intentions to complete The Bottom.
  • Task 2: Rotate the car’s bottom to assist the human in screwing in The Bottom.
  • Task 3: Collaborate with the human using the hierarchical framework to complete The Middle.
@@ V-C Results
V-C1 H1 - Time Efficiency
Table I compares the average completion times of three different frameworks—each utilizing different modalities—across the three sub-tasks outlined in the user study section V-B3. The table summarizes results from 40 trials for each task and modality. The percentage values in the table are calculated relative to the longest completion time recorded for each task. It is important to note that in Task 3, human intention cannot be accurately predicted using vision alone, as human trajectories exhibit minimal movement during the screwing process. This limitation prevents the intention prediction model from making accurate predictions based solely on visual input, resulting in an inability to complete Task 3 with the vision-only framework. Thus, an infinite completion time is assigned for this scenario. As the results show, the multimodal framework achieves the shortest completion time across all tasks compared to the single-modal frameworks. This demonstrates that the proposed framework effectively enhances the efficiency of human-robot collaboration.
Vision Audio Multimodal
Task1 (%) $75.70\pm 10.56$ $72.02\pm 9.10$ $\textbf{67.42}\pm\textbf{4.91}$
Task2 (%) $48.42\pm 20.90$ $43.72\pm 13.50$ $\textbf{39.45}\pm\textbf{14.83}$
Task3 (%) $\infty$ $74.19\pm 10.83$ $\textbf{65.83}\pm\textbf{3.86}$
TABLE I: The average completion time with standard deviation.
@@ V-C2 H2 - Disturbance Minimization
The presence of multiple humans in the same environment can introduce errors in human detection. As discussed in section IV-B, we propose a hierarchical human detection module to address the challenges and disturbances caused by multiple humans. For comparison, sample results from a naive human pose detection model (without hierarchical detection) are shown in fig. 6. As illustrated, the naive detection model can result in errors or misdetections when multiple humans are present. In contrast, the proposed hierarchical framework reduces these disturbances by prioritizing the human closest to and interacting with the robot, as demonstrated in fig. 7.
@@
Figure 7: Correct detection using the hierarchical human detection framework.
Hierarchical Human Detection. We conducted the following case study to numerically evaluate the effectiveness of the hierarchical framework. In this setup, one person stood close to the robot, serving as the human operator, while another person walked through the environment to introduce noise for the detection model. To evaluate the framework’s robustness, we measured the keypoint deviation over time. Figure 8 shows the keypoint deviation for three methods: naive detection without hierarchical human detection, naive detection with Kalman filtering, and hierarchical human detection. The ground truth is represented by a constant zero line (green dotted line). As illustrated, the naive detection method, even when smoothed with Kalman filtering, results in sharp fluctuations in the detected keypoints, with a variance of 0.013. In contrast, the proposed hierarchical human detection method (black curve) provides robust and stable detection, closely aligning with the ground truth. These results support the hypothesis that the hierarchical framework effectively minimizes disturbances in multi-human environments.
@@
Action prediction. Table VI compares the action prediction accuracy of the proposed hierarchical human detection framework against a naive baseline and four baselines using filtering methods. The methods are: (a) ”w/o hierar,” which predicts actions using a naive human pose; (b) ”median/wiener/kalman/ema filter,” which applies respective filters to the naive human pose before predicting actions; and (c) ”w/ hierar,” which uses the hierarchical detection framework for action prediction. In the table, ”Action1” corresponds to ”get_connectors,” ”Action2” to ”get_screws,” and ”Action3” to ”get_wheels.” ”No Action” indicates instances where the human’s action did not match any of these three actions. ”All” represents the overall average accuracy across all actions. The results clearly show that hierarchical human detection significantly improves action prediction accuracy compared to other methods.
@@
TABLE II: Comparison of action prediction accuracy.
@@
To minimize critical class-dependent misclassification errors—such as mistakenly classifying ”stay constant” as ”reach out to human”—we implement sensitivity-aware intention prediction. This approach strategically ”transfers” high-sensitivity errors to lower-sensitivity ones, improving overall prediction accuracy. Further details are available in section -C.
V-C3 H3 - Task Completion
As discussed in section IV-A, a multimodal observation framework significantly enhances the information captured from the environment. This approach reduces errors compared to single-modality perception systems, which often lack adequate information about humans and the environment. We evaluated the effectiveness of multimodal observations during the user study. Table III shows the task completion rates for three tasks in the user study, conducted under three different observation modalities. Note that Task 3 cannot be completed using a vision-only framework, as explained in section V-C1. As shown intable III, the multimodal framework allows the robot to respond more proactively by utilizing the richer information it gathers. This enables the robot to complete tasks more effectively and better meet user expectations.
Vision Audio Multimodal
Task1 (%) $70.63\pm 18.65$ $90.63\pm{6.75}$ $\textbf{93.75}\pm\textbf{9.32}$
Task2 (%) $68.75\pm 22.24$ $63.75\pm 18.35$ $\textbf{85.38}\pm\textbf{13.58}$
Task3 (%) 00 $81.25\pm 19.98$ $\textbf{96.25}\pm\textbf{4.37}$
TABLE III: Task completion rates under different modalities.
V-C4 H4 - Adaptibility

### v2

@@ IV-C Human Physical Action Prediction
Rather than predicting human actions solely based on current pose trajectories, our approach first predicts future trajectories and then infers actions using both the current and predicted trajectories. This can enhance system performance by ensuring that the action prediction is informed by anticipated future movements [4]. We jointly train trajectory and action prediction models using supervised learning, utilizing collected offline data. The trajectory prediction model is optimized using regression loss (L2 loss), while the action prediction model employs classification loss (cross-entropy loss). Once trained, the action prediction model $\hat{f}_{\text{action}}$ predict human actions $\hat{a}_{H_{p}}^{t}$ and its output confidence based on the current human pose: $\hat{a}_{H_{p}}^{t},\hat{\epsilon}_{H_{p}}^{t}=\hat{f}_{\text{action}}(\hat{z}_{H}^{t}).$
Error Type Sensitivity score
“No Action” predicted as other actions high
One action predicted as another action medium
One action predicted as ”No Action” low
TABLE I: Sensitivity score of different misclassification errors.
Additionally, we assign different sensitivity scores to various types of misclassification errors and optimize the prediction model to minimize the most critical errors. The sensitivity scores assigned to each type of error are shown in Table. I. We propose two restriction strategies to mitigate higher-sensitive errors: confidence and boundary restriction, the details and experimental results are shown in Sec. V-C2.
@@
To enhance model generalization to different human behaviors, we implement an online adaptation method [2] during deployment, enabling the robot to refine predictions in real-time by learning each user’s behavior [11]. Our joint training framework supports dynamic adaptation by optimizing the trajectory prediction model online using observed trajectory feedback. As action predictions depend on the trajectory module, updates to the trajectory model directly improve action prediction accuracy [1]. The online adaptation approach implements the method reported in [3]. This technique significantly improves prediction performance, ensuring that the robot’s responses accurately align with each user’s behaviors and preferences.
IV-D Human Speech Action Recognition
For the audio modality $O_{A}$, the robot recognizes speech action $\hat{a}_{H_{s}}^{t}$ along with its confidence using a pretrained speech recognition model $\hat{f}_{\rm speech}$: $\hat{a}_{H_{s}}^{t},\hat{\epsilon}_{H_{s}}^{t}=\hat{f}_{\rm speech}(O_{A})$. In deployment, this model’s parameters are fixed. To enhance the quality of speech recognition, a Voice Activity Detector (VAD) is utilized to filter the input. This VAD activates the speech recognition module only when significant audio signal energy, not noise, is detected.
IV-E Hierarchical Human Plan Prediction
After obtaining human physical action $\hat{a}_{H_{p}}^{t}$ and human speech action $\hat{a}_{H_{s}}^{t}$, the robot determine the final human action $\hat{a}_{H}^{t}$ based on priority (speech command “stop” has the highest priority. In other cases, priority is determined by the action with the higher confidence level). The robot then predicts human plan $\hat{x}_{H}^{t}$ based on the final human action $\hat{a}_{H}^{t}$. We divide this process into two hierarchical levels. Low-level plan prediction focuses on short-term sub-task nodes, such as “get screws”, which dictate immediate action sequences. High-level plan prediction captures broader human intent, corresponding to long-term task sequences.
@@ IV-E Hierarchical Human Plan Prediction
The robot identifies the reference node sequence $R^{\ast}$ within the task graph that most closely aligns with the recorded task progress $P^{t}$. This alignment is formulated as:
@@
where $d$ represents the distance between the observed task progress $P^{t}$ and each possible reference sequence $R$ within the task graph. The distance metric $d$ is calculated using the dynamic time warping (DTW) algorithm [31]. The result, $R^{\ast}$, represents the best-matched reference node sequence. Note that $R^{\ast}$ may not be unique, as multiple node sequences may present equal distances, allowing for different pathways to the same goal. For instance, in the long-term task graph Fig. 1, the assembly process may start from node 1, 3, or 5, indicating different initial conditions. Suppose the observed task sequence suggests that the human is assembling the short tubes (Plan 3). The robot aligns the observed task progress $P^{t}$ with the closest reference sequence $R^{\ast}$ within the plan graph.
Following this alignment, the human plan prediction is obtained by refining the predicted human action based on the most probable action class within $R^{\ast}$. This adjustment is formulated as:
$\hat{x}_{H}^{t}=R^{\ast}(\hat{a}_{H}^{t}),$ (5)
where $\hat{x}_{H}^{t}$ is the predicted human plan, aligning $\hat{a}_{H}^{t}$ with the most likely action from $R^{\ast}$. For example, as shown in the sub-task graph of Fig. 1, in Plan 1 of assembling tubes and connectors, if the human has already performed action A (fetching connectors), the most probable next action in $R^{\ast}$ is Action B (fetching tubes and assembling them). The robot adjusts its prediction accordingly, ensuring its motion planning remains synchronized with human execution.
By hierarchically structuring human plan prediction into two levels and aligning the robot’s predictions with the task graph and task progress, this method improves the accuracy and flexibility of human plan prediction.
IV-F Robot Motion Controller
Upon acquiring the human plan $\hat{x}_{H}^{t}$, the robot’s motion planner generates the robot motions according to human plan and current task progress:
$q^{t:t+m}$ $=M_{R}(\hat{x}_{H}^{t})$ (6)
where $q^{t:t+m}$ represent the robot motion sequences and $M_{R}$ represents the motion controller. The motion controller operates as a feedback controller with an update frequency of approximately 30 Hz. In accordance with the optimization objective, accurate prediction of $\hat{x}_{H}^{t}$ is crucial for optimal motion generation within the task graph. Errors in $\hat{x}_{H}^{t}$ disrupt robot motions, reducing efficiency and robustness in HRC.
@@ V-A Toy Car Assembly Task
We evaluate the effectiveness of our proposed framework through a long-term HRC task, specifically the toy car assembly task, as shown in Fig. 3. This assembly process incorporates various components, including both short and long tubes, connectors, screws and wheels. This task is divided into four stages: assembling the bottom, the middle, the top of the car, and installing the wheels, as shown in Fig. 1. For simplicity, we will refer to the four stages as The Bottom, The Middle, The Top and Install Wheels. The assembly order for The Middle and The Top is flexible, allowing variations in the sequence order.
Figure 3: Final Assembly Task
Each stage is designed to require collaborative efforts, in which the robot’s role is to support the human collaborator by delivering necessary objects and performing specific assistive actions, such as rotating or lifting objects. The actions and motion sequences within these stages are non-deterministic, offering a range of valid action sequences that can successfully complete each sub-task. This flexibility means that the specific sequence of actions may vary according to individual human preferences. For instance, in Fig. 1, completing Node 1 (or Node 5), which involves assembling tubes and connectors, can be achieved through various action sequences depending on the human’s plan. Consider Plan 2 as an example: the human begins by getting the connectors (Node A). At this point, the robot predicts the human’s plan, based on a default prediction model, that the human intends to ”get the short tube” (Node E). However, the human instead verbalizes ”long” (Node D), correcting the robot’s prediction, which then adjusts its action to get the long tube (Node F). After obtaining the correct component, the human proceeds to assemble the tube (Action B). Subsequently, the human gets the connectors again, and this time, the robot correctly predicts the human’s intention to ”get the short tube,” aligning with the expected plan. Through dynamic adjustments to its assistance based on predicted human actions, the robot ensures seamless and flexible collaboration across various human plans, ultimately leading to the efficient completion of the overall task.
Figure 4: Environment Setting
The environment setup is shown in Fig. 4, as the tools and objects needed for the assembly task are positioned across different locations. Connectors, screws, wheels, and screwdrivers are placed near the human to allow easy access. Conversely, the tubes are stored on a table that is closer to the robot arm and relatively farther from the human. The locations of all objects are fixed, and we assume that both the robot and the human know all the exact locations and there is no uncertainty in object perception. During the assembly process, the robot assists according to its understanding of human plans, current task progress and current environment. For instance, if the human plans to assemble tubes by reaching for connectors, the robot responds by delivering the appropriate tube—either short or long—based on the observed signals. When the human plans to retrieve screws, the robot aids by spinning the base-object to facilitate easier handling. Similarly, when the task involves attaching wheels, the robot assists by lifting the car’s body to better position it for wheel installation. The source code and demonstration video for the HRC toy car assembly task are available at https://github.com/intelligent-control-lab/Robust-Hierarchial-Multimodal-HRC.
@@ V-B1 Implementation Details
  • Human Distance Estimation. We first use a pretrained MobileNet model [29] to detect “human” from RGB data, applying a confidence threshold of 0.6. Then we estimate human distance from corresponding depth image.
  • Human Pose Estimation. The pretrained BlazePose model [8] extracts 33 key points, of which we focus on 15 upper-body points for physical action prediction.
  • Human Physical Action Prediction. We implement the DLinear model [35] for action and trajectory prediction, which decomposes input time series into trend and seasonal components. This model is trained using offline data to be specified in Sec. V-B4. We predict four human physical actions: “get connectors”(Action1), “get screws”(Action2), “get wheels”(Action3), and No Action, where “No Action” indicates instances where the human’s action did not match any of other three actions.
  • Human Speech Action Recognition. The DeepSpeech model [7], an open-source Speech-to-Text engine is utilized for recognizing and processing speech commands. We recognize five speech commands: “short”, “long”, “spin”, “lift”, and “stop”.
@@ V-B2 Hypothesis
  • H1: The framework increases task success rates in collaborative tasks.
@@
  • H3: The framework reduces task completion time and enhances the efficiency of HRC.
@@
  • H5: The framework allows users to be more flexible and improves user satisfaction in collaboration tasks.
@@ V-B3 User Study
To evaluate the proposed framework and test the five hypotheses, we conducted a user study involving 10 participants. The study divides the toy car assembly task into three consecutive sub-tasks:
  • Task 1: Assemble tubes and connectors in The Bottom Stage.
  • Task 2: Based on the completion of The Bottom Stage, complete The Middle Stage and The Top Stage according to the user plan.
  • Task 3: Based on the completion of the previous three stages, complete Install Wheels Stage.
We define “Perception Modality = 0” (referred to “PM” below) as vision-only perception, “PM = 1” as audio-only perception and “PM = 2” as multimodal perception, “Plan Prediction = 0” (referred to as “PP” below) as having no plan prediction and “PP = 1” as having plan prediction. When “PP = 0”, the robot has a pre-defined plan sequence and only reacts after the human takes certain actions. When “PP = 1”, the robot hierarchically predicts human plans and reacts proactively. By defining these two variables, we have six groups of test conditions. Under each group, every user performs each task using any plan for four times. Thus, there will be 120 trials in each group and we collect 720 trials in total.
@@ V-C Results
V-C1 H1 - Task Success Rate
Table. II compares the task success rates of six groups for each of the three sub-tasks outlined in the user study in Sec. V-B3. The table summarizes results from 40 trials for each task and group. As shown in Table. II, the group with multimodal perception and hierarchical plan prediction (“PP = 1, PM = 2”) achieves the highest task success rate (91.79%). Notably, Task 2 demonstrates the most pronounced benefit of multimodal perception (“PM = 2”). In this task, the vision-only group (“PM = 0”) struggles significantly when attempting to prioritize the Middle Stage, resulting in a big decline of success rate (40.5% on average compared to “PM = 2”). Furthermore, integrating hierarchical plan prediction with multimodal perception (“PP = 1, PM = 2”) increases success rate by 14.22% compared to without plan prediction (“PP = 0, PM = 2”). These results demonstrate the synergistic advantages of combining hierarchical plan prediction and multimodal perception, particularly in handling complex tasks and user plans.
PP PM Task1 (%) Task2 (%) Task3 (%)
0 0 $74.45\pm 17.15$ $48.34\pm 6.21$ $56.46\pm 18.38$
1 $92.34\pm 4.68$ $82.32\pm 19.18$ $69.58\pm 16.74$
2 $89.12\pm 9.56$ $83.21\pm 12.21$ $60.45\pm 17.39$
1 0 $70.63\pm 18.65$ $50.12\pm 8.26$ $68.75\pm 22.24$
1 $90.63\pm 6.75$ $81.25\pm 19.98$ $63.75\pm 18.35$
2 $\textbf{93.75}\pm\textbf{9.32}$ $\textbf{96.25}\pm\textbf{4.37}$ $\textbf{85.38}\pm\textbf{13.58}$
TABLE II: Comparison of task success rates.
@@ V-C2 H2 - Disturbance Minimization
The presence of multiple humans in the same environment can introduce errors in human detection. As discussed in Sec. IV-B, we propose a hierarchical human detection module to address the challenges and disturbances caused by multiple humans. For comparison, sample results from a naive human pose detection model (without hierarchical detection) are shown in Fig. 5. As illustrated, the naive detection model can result in errors or misdetections when multiple humans are present. In contrast, the proposed hierarchical framework reduces these disturbances by prioritizing the human closest to and interacting with the robot, as demonstrated in Fig. 6.
@@
Figure 6: Correct detection using the hierarchical pose detection framework.
Human Keypoints Detection. We conducted the following case study to numerically evaluate the effectiveness of the hierarchical pose detection method on human keypoints detection. In this setup, one person stood close to the robot, serving as the human operator, while another person walked through the environment to introduce noise for the detection model. To evaluate the method’s robustness, we measured the keypoint deviation over time. Fig. 7 shows the keypoint deviation for three methods: naive detection without hierarchical human detection, naive detection with Kalman filtering, and hierarchical human detection. The ground truth is represented by a constant zero line (green dotted line). As illustrated, the naive detection method, even when smoothed with Kalman filtering, results in sharp fluctuations in the detected keypoints, with a variance of 0.013. In contrast, the proposed hierarchical human detection method (black curve) provides robust and stable detection, closely aligning with the ground truth. These results support the hypothesis that the hierarchical pose detection method effectively minimizes disturbances in multi-human environments.
@@
Action prediction. Table. III compares the action prediction accuracy of the proposed human physical action prediction module against a naive baseline and four baselines using filtering methods in multi-human noisy environments. The methods are: (a) “w/o hierar,” which predicts actions using a naive human pose; (b) “median/wiener/kalman/ema filter,” which applies respective filters to the naive human pose before predicting actions; and (c) “w/ hierar,” which uses the hierarchical detection framework for action prediction. In the table, “All” represents the overall average accuracy across all actions. The results clearly show that hierarchical human detection significantly improves action prediction accuracy compared to other methods.
@@
TABLE III: Comparison of action prediction accuracy.
@@
To minimize critical class-dependent misclassification errors—such as mistakenly classifying “stay constant” as “reach out to human”—we implement sensitivity-aware action prediction, which strategically ”transfers” higher-sensitivity errors to low-sensitivity ones as defined in Sec. V-B1, improving overall prediction accuracy. We propose two restriction strategies and conduct experiments on offline data to demonstrate their effectiveness.
The first strategy is confidence-based restriction, which assign different confidence scores to different actions. If the predicted probability for an action $a_{1}$ (other than ”no-action”) is lower than a confidence threshold (i.e 0.6), we label the action as ”no action” $a_{0}$ instead, even if the original probability of the action $a_{1}$ is higher than ”no action”. This approach helps ”transfer” a high-sensitivity error to a lower-sensitivity error.
The second strategy is the physical ”boundary” restriction. The ”boundary” is a specific range of working area where a human operates when doing specific actions. We consider both the current and predicted human trajectory to determine whether the human’s hand is within the working area boundary. If the hand satisfied boundary restriction, we proceed with the action prediction model. Otherwise, we label the action as ”no action.” The error rates for different sensitivities under various strategies are shown in Fig. 8. As demonstrated, the combination of ”confidence + boundary” restriction effectively reduces the two types of most sensitive error rates.
Figure 8: Action prediction error rate under different restriction strategies
These results demonstrate that the hierarchical human pose detection method and the sensitivity-aware method action prediction method used in our HRC framework significantly mitigate the environment disturbances and improve the action prediction accuracy.
V-C3 H3 - Time Efficiency
Similar to Sec. V-C1, Table. IV compares the average completion times of six experimental groups across three sub-tasks, expressed as percentages relative to the longest recorded time for each task. As the results show, the group with multimodal perception and hierarchical plan prediction (”PP = 1, PM = 2”) achieves the shortest completion time across all tasks, with an average reduction of 15.85% compared to groups lacking both plan prediction and multimodal perception (”PP =0, PM = 0&1”). Specifically, group with multimodal perception (”PM = 2”) reduces task completion time by 8.54% compared to single-modality groups (”PM = 0&1”), while hierarchical plan prediction (”PP = 1”) further decrease completion times by 7.42% compared to groups without it (”PP = 0”). These results demonstrate that the proposed framework effectively combines the capabilities of multimodal perception and hierarchical plan prediction to optimize task efficiency, enabling seamless and proactive human-robot collaboration.
PP PM Task1 (%) Task2 (%) Task3 (%)
0 0 $78.80\pm 13.57$ $86.80\pm 10.56$ $53.19\pm 18.31$
1 $80.05\pm 7.23$ $79.34\pm 12.49$ $62.33\pm 14.54$
2 $75.86\pm 15.67$ $72.56\pm 6.12$ $45.12\pm 13.29$
1 0 $75.70\pm 10.56$ $80.46\pm 13.98$ $48.42\pm 20.90$
1 $72.02\pm 9.10$ $74.19\pm 10.83$ $43.72\pm 13.50$
2 $\textbf{67.42}\pm\textbf{4.91}$ $\textbf{65.83}\pm\textbf{3.86}$ $\textbf{39.45}\pm\textbf{14.83}$
TABLE IV: The average completion time with standard deviation (only calculate success cases).
V-C4 H4 - Adaptability